OpenStudy (anonymous):

Express as a single logarithm: log(x^2+6x+8)-log(x^2-16)

OpenStudy (anonymous):

Do you know the log rule:\[\ln (a)-\ln(b) = \ln(\frac{ a }{ b })\]

OpenStudy (anonymous):

im learning it, I am a little uncertain as to how to set it up.

OpenStudy (mathstudent55):

Here's a simple example of how to apply the rule above: \(\log (x^2)-\log(x) = \log(\dfrac{ x^2 }{ x }) = \log x\) Now apply the rule to your problem. Then factor the numerator and denominator and reduce the fraction.

OpenStudy (anonymous):

x+2/x?

OpenStudy (mathstudent55):

What happened to "log"?

OpenStudy (anonymous):

Not quite, its (x+2) for the numerator, on the top, but its not x for the denominator. how did you start?

OpenStudy (anonymous):

log=(x+2)/(x-4)?

OpenStudy (mathstudent55):

\(\log(x^2+6x+8)-\log(x^2-16)\) \(= \log \dfrac{x^2+6x+8}{x^2-16}\) Now factor the numerator and denominator. What do you get?

OpenStudy (anonymous):

num i got (x+4)(x+2)

OpenStudy (mathstudent55):

\(= \log \dfrac{x^2+6x+8}{x^2-16}\) \(= \log \dfrac{(x + 4)(x + 2)}{(x+4)(x-4)}\) \(= \log \dfrac{\cancel{(x + 4)}(x + 2)}{\cancel{(x+4)}(x-4)}\) \(= \log \dfrac{x + 2}{x - 4}\)

OpenStudy (anonymous):

Thanks man you ROCK! I understand clearly now

OpenStudy (mathstudent55):

You're welcome.